OpenStudy (anonymous):

When solving x2 - 20x = 5 by completing the square what value is added to both sides of the equation? -4 4 50 100?

OpenStudy (ash2326):

We have \[x^2-20x=5\] We need to get the form \[(x-a)^2=x^2-2ax+a^2\] here 2a=20 so a=10 we need to add 100 both sides \[x^2-20x+100=105\] so we get \[(x-10)^2=105\]